OpenStudy (anonymous):

Let f(x) = x – 6 and g(x) = x2 – x + 4. Find g(f(x)). x2 – 13x + 46 x2 – 11x + 34 x2 – 2x – 2 x2 – x – 2

OpenStudy (anonymous):

f(x) = x - 6 g(x) = x^2 - x + 4 g(f(x)) = (x - 6)^2 - (x - 6) + 4 = x^2 - 12x + 36 - x + 6 + 4 = x^2 - 13x + 46